OpenStudy (anonymous):

Express the sum in the simplest form. (3x/x+5) + (7/x+5)

OpenStudy (anonymous):

common denominators so add the numerators then copy the denominators \[\frac{3x}{x+5} + \frac{7}{x+5} = \frac{3x + 7}{x+5}\]